OpenStudy (anonymous):

What is the distance from (−3, 1) to (−1, 5)? Round your answer to the nearest hundredth.

OpenStudy (igreen):

Formula for the distance between two points: \(\sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\) \(x_2 = -1\) \(x_1 = -3\) \(y_2 = 5\) \(y_1 = 1\)

OpenStudy (igreen):

Can you plug them in and solve? @Bballplayer48

OpenStudy (tkhunny):

Have you considered the Distance Formula? iGreen did.

OpenStudy (anonymous):

No I don't know it

OpenStudy (anonymous):

1) 7.9 2) x = 7 3) y = 4

OpenStudy (anonymous):

@iGreen Do I find the square root? at the end

OpenStudy (tkhunny):

Well, there you go, then. Use the substitution indicated by iGreen and you'll be done in no time.

OpenStudy (anonymous):

Ok ty

OpenStudy (anonymous):

1/ Distance Formula: Given the two points (x1, y1) and (x2, y2), the distance between these points is given by the formula: sqrt ( (x2-x1)² + ( y2-y1) ) 2 and 3/ the Midpoint Formula is : ( (x1+x2)/2, (y1+y2)/2 )

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

problem solved

OpenStudy (igreen):

@Bballplayer48 Yes, you do..

OpenStudy (igreen):

Tell me what you get. @Bballplayer48

OpenStudy (anonymous):

I got 2.449? @iGreen

OpenStudy (igreen):

I don't think that's correct.

OpenStudy (igreen):

Can you tell me how you got it?

OpenStudy (anonymous):

It isn't

OpenStudy (anonymous):

3.60?

OpenStudy (igreen):

\(\sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\) \(\sqrt{((-3)-(-1))^2 + (5-1)^2}\) Subtract -3 - -1, what do you get?

OpenStudy (anonymous):

OHHHH I forgot to put the exponent up top

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

@iGreen I got 8

OpenStudy (anonymous):

But thats wrong to :(

OpenStudy (anonymous):

Forget it :(

OpenStudy (igreen):

-3 - -1 = -2 (-2)^2 = 4 5 - 1 = 4 4^2 = 16 16 + 4 = 20 sqrt of 20 = 4.47

OpenStudy (igreen):

@Bballplayer48